OpenStudy (anonymous):

substitution method 5x-2y=-15 2x-y=-5

OpenStudy (anonymous):

You must get x or y by itself in one of the equations... I choose y in the second equation. Add 5 and add y to each side. You now have 2x+5=y Substitute 2x+5 for y in the first equation. 5x-2(2x+5)=-15 Simplify by distributing the 2. 5x-4x-10=-15 Get x by itself now. x-10=-15 Add 10 to each side. x=-5 Now substitute -5 for x in either equation. I choose the second again because it is simpler. 2(-5)-y=-5 -10-y=-5 Add 5 and add y to both sides. -5=y Your coordinates are (-5, -5)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Bethany your answer is incorrect. Try placing your x and y into one of the original equations... It doesn't come out with the correct answer.

OpenStudy (blurbendy):

brookew4a is correct

OpenStudy (anonymous):

Thanks blur :)

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

for emlimination method 3x+2y=6 6x+4y=-3 i got 36=9 is that correct